Use the compound interest formulas A = p( 1+r/n)^nt and A =pe^rt to Solve.
Find the accumulated value of an investment of $1130 at 8% compounded annually for 14 years.

Answer by mathmate(429)

Question:
Use the compound interest formulas A = p( 1+r/n)^nt and A =pe^rt to Solve.
Find the accumulated value of an investment of $1130 at 8% compounded annually for 14 years.

Solution:
The formulas are correct but not applicable to the present case.
We use
Future value with yearly investment A for n years at interest i
F=A/(i)
=1130/(.08)
=27362.86